# An early dark matter-dominated phase in the assembly history of Milky Way-mass galaxies suggested by the TNG50 simulation and JWST observations

### v1

@@ 1 Introduction
Constraints on the mass budget, i.e. the relative contribution of dark matter, gas and stellar mass, can offer insight into the mass assembly histories of galaxies. In the local Universe, moderately massive galaxies ($M_{*}\sim 10^{10-11}\,\rm M_{\odot}$) are baryon-dominated near their centers ($\sim 1\,\rm kpc$), but contain substantial dark matter mass within the effective radius ($f_{\rm DM}(<r_{\rm e})\sim 40-70\%$ for $r_{\rm e}\sim 5-10\,\rm kpc$; Martinsson et al. 2013). Studies of ionized gas kinematics at $z\sim 1-3$ paint a very different picture, as these find an increase in the baryonic mass fraction within $r_{\rm e}$ toward higher redshift for galaxies of $M_{*}\gtrsim 10^{9.5}\,\rm M_{\odot}$, with galaxies appearing almost entirely baryon-dominated at $z\sim 2$ (Wuyts et al., 2016; Price et al., 2016, 2020; Genzel et al., 2017, 2020; Nestor Shachar et al., 2023). For very massive galaxies at $z\sim 2$ Genzel et al. (2017, 2020) propose this may imply that the dark matter distributions are cored instead of cuspy, which may be caused by a very rapid formation process coupled with stellar and black hole feedback.
The James Webb Space Telescope (JWST) enables the measurement of galaxy kinematics at even higher redshifts and also lower masses, thereby probing the possible progenitors of more massive galaxies at cosmic noon. In de Graaff et al. (2023) we used JWST/NIRSpec observations to measure the ionized gas kinematics for six small ($r_{\rm e}\sim 1\,\rm kpc$) galaxies at $z\sim 6-7$ of low stellar mass, $M_{*}\sim 10^{7-9}\,\rm M_{\odot}$. Interestingly, the dynamical masses inferred from the ionised gas kinematics exceed the stellar masses by a factor $\sim 30$, and the baryonic masses by a factor $\sim 3$, which may imply surprisingly high dark matter fractions of $f_{\rm DM}(<r_{\rm e})\approx 0.7$. However, these JWST results are still subject to large observational and systematic uncertainties, raising the question whether these implied dark matter fractions are realistic, and whether these are compatible with observations at intermediate redshifts.
@@ 2 TNG simulations
[…]
@@ 3 Mass budget at $z\sim 6$
We start assessing the central mass budget of high-redshift galaxies by looking at our sample of TNG50 central galaxies at $z=6$, which is the median redshift of the JWST measurements by de Graaff et al. 2023, and measuring the baryonic mass fraction in a (3D) spherical aperture of the stellar half-mass radius:
@@
Figure 1 shows $f_{\rm baryon}(<r_{\rm e})$ versus the total stellar mass in TNG50, color-coded by $r_{\rm e}$: there is a strong dependence of $f_{\rm baryon}$ on stellar mass. Galaxies of low stellar mass ($M_{*}\lesssim 10^{8.5}\,\rm M_{\odot}$) have baryon fractions in the range $f_{\rm baryon}(<r_{\rm e})\sim 0.1-0.3$ and thus are dark matter-dominated. Above $M_{*}\sim 10^{8.5}\,\rm M_{\odot}$, $f_{\rm baryon}(<r_{\rm e})$ increases rapidly in TNG50, and the most massive systems ($M_{*}\sim 10^{10}\,\rm M_{\odot}$) are baryon-dominated. Moreover, $f_{\rm baryon}(<r_{\rm e})$ is covariant with $r_{\rm e}$ at fixed stellar mass, so that larger effective radii imply lower values of $f_{\rm baryon}$.
@@
These simulation results are in remarkably good agreement with the measurements from de Graaff et al. (2023), inferred from JWST observations of low-mass galaxies ($M_{*}\sim 10^{7-9}\,\rm M_{\odot}$) at $z\sim 6-7$. These observational estimates of the baryonic mass were obtained from spectral energy distribution (SED) modeling to low-resolution spectroscopy, while the total mass estimates were inferred dynamically from measurements of ionized gas kinematics. We emphasize that these observed estimates differ conceptually in important ways from those for the simulated baryon fractions. For example, the observed effective radius represents the projected half-light radius of the emission line instead of a half-mass radius, although we find that these are $\sim 1\,\rm kpc$ in both simulated and observed galaxies. The stellar masses were inferred from SED modeling and the cold-gas mass is empirically inferred from the star formation rate (SFR); both observational estimates carry significant systematic uncertainties. Moreover, observations are subject to projection effects that can lead to an underestimated circular velocity. This may for example explain the data point for which the baryonic mass exceeds the total dynamical mass ($M_{\rm baryon}>M_{\rm dyn}$). In light of these different issues, the broad agreement found between the observations and the TNG50 data shown in Figure 1 is striking.
@@
Figure 1: Baryonic mass fraction within the stellar half-mass radius (TNG50) and the emission line half-light radius (JWST; de Graaff et al. 2023) of galaxies vs. total stellar mass. Throughout, by baryonic fraction we mean the ratio between the mass in gas and stars and the total mass of the galaxy (see Eq. 1). The TNG50 simulation predicts low baryon fractions for $M_{*}<10^{9}\,\rm M_{\odot}$ and a sharp increase in $f_{\rm baryon}$ toward higher masses. The half-mass radii are small ($\sim 1\,\rm kpc$ in physical units) and, at fixed stellar mass, are anti-correlated with the baryon fraction. With the exception of one source, which may have high $f_{\rm baryon}$ due to an underestimated dynamical mass, the baryon fractions measured at $z\sim 6$ with JWST are in good agreement with the simulated galaxies.
@@ 4.2 The central mass component budget:
[…]
Turning to the descendants of the selected $z=6$ galaxies in TNG50, we find that for any individual descendant $f_{\rm baryon}(<1\,\rm kpc)$ increases rapidly toward lower redshift: while the central region is still dark matter-dominated at $z=6$ with $f_{\rm baryon}(<1\,\rm kpc)\sim 0.25$, it is highly baryon-dominated at $z\sim 2$ $(f_{\rm baryon}(<1\,\rm kpc)\sim 0.7$). Considering the strong dependence of $f_{\rm baryon}$ on $M_{*}$, this evolution can be understood by the rapid growth of total stellar mass over time found in Figure 2.
@@
Toward lower redshift the stellar-to-gas mass ratio increases sharply for the $z=6$ descendants, and this central build-up of stellar mass coincides with the rapid growth of the total stellar mass (Figure 2). At $z\lesssim 2$ we find that $M_{\rm gas}(<1\,\rm kpc)\sim 0$, and the mass ratio therefore plateaus at $M_{*}/M_{\rm gas}(\lesssim 1\,\rm kpc)\sim 100$ due to our imposed gas mass floor. This likely reflects the kinetic AGN feedback mode in the IllustrisTNG model (Pillepich et al., 2018b), in which black hole-driven winds remove gas from their surroundings.
@@ 4.3 Comparison with observations at cosmic noon
In Figure 4 we show the dark matter fraction, $f_{\rm DM}(<r_{\rm e})\equiv 1-f_{\rm baryon}(<r_{\rm e})$, as a function of redshift. For the TNG50 $z=6$ descendant sample the median evolution in $f_{\rm DM}$ (solid line) is color-coded by the median stellar mass at the given redshift, and contours indicate the variation within the population. We also show results at $z\sim 1-3$ from various ground-based studies that obtained dynamical masses from ionized gas kinematics (Wuyts et al., 2016; Genzel et al., 2020; Price et al., 2020; Puglisi et al., 2023). All observed data points are color-coded by the total stellar mass, typically inferred from SED modeling to photometric data.
@@ 5 Discussion & Conclusions
We have used the cosmological simulation TNG50 to extract theoretical expectations for the stellar, gas and dark matter mass components in the central parts of high-redshift galaxies. We have explored these components’ fractional contributions as a function of stellar mass and redshift. In particular, we have shown that TNG50 predicts that the central baryon fraction of galaxies is independent of redshift at fixed stellar mass: the centers of low stellar mass galaxies have always been dark-matter dominated. The central baryon fraction, is however, a strong function of stellar mass at all redshifts, and so greatly increases along the evolutionary paths of individual galaxies. The central stellar-to-gas mass ratio, on the other hand, grows strongly towards lower redshifts, even at a fixed stellar mass.
Turning to high redshifts, we compare these results to recent measurements of the mass budget at $z\sim 6$ from JWST observations of galaxies of low stellar mass ($M_{*}\sim 10^{7-9}\,\rm M_{\odot}$; de Graaff et al., 2023). Overall, these observational results appear in good agreement with TNG50 galaxies at $z=6$: the baryonic mass fractions are approximately equally low at low stellar masses $f_{\rm baryon}(<1\,\rm kpc;M_{*}=10^{8-9}\,\rm M_{\odot})\sim 0.2-0.4$. The relative mass fraction among the baryonic components (stars and gas) also agrees at least qualitatively. We find $M_{*}(<1\,\rm kpc)\ll M_{\rm gas}(<1kpc)$ for both the simulated and observed galaxies at $z\sim 6$ and $M_{*}<10^{9}\,\rm M_{\odot}$. Quantitatively, on the other hand, we find that – at face value – the stellar-to-gas mass ratio is a factor two higher in the TNG50 simulated systems compared to the observations. However, we deem this comparison inconclusive as this mass ratio in TNG depends on the resolution of the simulation (at the $20-40\%$ level across a factor of 16 in mass resolution at $z\sim 5-6$ and $M_{*}\sim 10^{8-9}\,\rm M_{\odot}$).
@@
The comparison in the mass budget between the high-redshift observations and simulations demonstrates the great potential constraining power of such observations. The TNG50 simulation makes clear predictions for the central baryon fraction and its dependence on stellar mass, and (lack thereof) on redshift. This is similarly true for the stellar-to-gas mass ratio, which also appears more sensitive to the details of the simulation (resolution, subgrid model). Therefore, systematically mapping these mass ratios as a function of stellar mass and redshift through observations would provide direct insight into galaxy mass assembly. Combined with a careful modeling analysis of observational biases by the use of mock observations from simulations, these observations will be able to place strong constraints on theoretical models and constrain the fundamental question: how efficient was star formation in the early universe?

### v2

@@ 1 Introduction
Constraints on the mass budget, i.e. the relative contribution of dark matter, gas and stellar mass, can offer insight into the mass assembly histories of galaxies. In the local Universe, moderately massive galaxies ($M_{*}\sim 10^{10-11}\,\rm M_{\odot}$) are baryon-dominated near their centers ($\sim 1\,\rm kpc$), but contain substantial dark matter mass within the effective radius ($f_{\rm DM}(<r_{\rm e})\sim 40-70\%$ for $r_{\rm e}\sim 5-10\,\rm kpc$; Martinsson et al. 2013). Studies of ionized gas kinematics at $z\sim 1-3$ paint a very different picture, as these find an increase in the baryonic mass fraction within $r_{\rm e}$ toward higher redshift for galaxies of $M_{*}\gtrsim 10^{9.5}\,\rm M_{\odot}$, with galaxies appearing almost entirely baryon-dominated at $z\sim 2$ (Wuyts et al., 2016; Price et al., 2016, 2020; Genzel et al., 2017, 2020; Nestor Shachar et al., 2023), although recent work suggests this may not be the case for all galaxies at this epoch (Sharma et al., 2023). For very massive galaxies at $z\sim 2$ Genzel et al. (2017, 2020) propose that such high baryon fractions may imply that the dark matter distributions are cored instead of cuspy, which may be caused by a very rapid formation process coupled with stellar and black hole feedback.
The James Webb Space Telescope (JWST) enables the measurement of galaxy kinematics at even higher redshifts and also lower masses, thereby probing the possible progenitors of more massive galaxies at cosmic noon. In de Graaff et al. (2024) we used JWST/NIRSpec observations to measure the ionized gas kinematics for six small ($r_{\rm e}\sim 1\,\rm kpc$) galaxies at $z\sim 6-7$ of low stellar mass, $M_{*}\sim 10^{7-9}\,\rm M_{\odot}$. Interestingly, the dynamical masses inferred from the ionized gas kinematics exceed the stellar masses by a factor $\sim 30$, and the baryonic masses by a factor $\sim 3$, which may imply surprisingly high dark matter fractions of $f_{\rm DM}(<r_{\rm e})\approx 0.7$. However, these JWST results are still subject to large observational and systematic uncertainties, raising the question whether these implied dark matter fractions are realistic, and whether these are compatible with observations at intermediate redshifts.
@@ 2 TNG simulations
We base our main analysis on the TNG50 simulation (Nelson et al., 2019a; Pillepich et al., 2019), the highest resolution simulation in the IllustrisTNG project (TNG hereafter; Marinacci et al., 2018; Naiman et al., 2018; Nelson et al., 2018; Pillepich et al., 2018a; Springel et al., 2017). TNG50 assumes a $\Lambda$CDM cosmology with cosmological parameters from the Planck Collaboration (2016), and uses the moving mesh code AREPO (Springel, 2010) and the IllustrisTNG galaxy formation model (Weinberger et al., 2017; Pillepich et al., 2018b) to simulate several thousand galaxies simultaneously within a cosmological volume of $\sim 50^{3}$ comoving Mpc3.
With a baryonic mass resolution of $m_{\rm baryon}=8.5\times 10^{4}\,\rm M_{\odot}$, a dark matter particle resolution of $m_{\rm DM}=4.5\times 10^{5}\,\rm M_{\odot}$ and a Plummer-equivalent gravitational softening length of $\epsilon^{z=0}_{\rm DM,*}=0.29\,\rm kpc$ at $z=0$, TNG50 has both the large volume and the high resolution needed to robustly evaluate the average mass profiles of galaxies on scales of $\sim 1\,\rm kpc$ across a wide range in redshift and stellar mass. To test the convergence of our results, we also use the TNG100 simulation (Nelson et al., 2019b), which employs the same galaxy formation model, but for a larger volume of $\sim 100^{3}$ comoving Mpc3 and lower resolution ($m_{\rm baryon}=1.4\times 10^{6}\,\rm M_{\odot}$; $m_{\rm DM}=7.5\times 10^{6}\,\rm M_{\odot}$; $\epsilon^{z=0}_{\rm DM,*}=0.74\,\rm kpc$).
@@ 3 Mass budget at $z\sim 6$
We start assessing the central mass budget of high-redshift galaxies by looking at our sample of TNG50 central galaxies at $z=6$, which is the median redshift of the JWST measurements by de Graaff et al. 2024, and measuring the baryonic mass fraction in a (3D) spherical aperture of the stellar half-mass radius:
@@
Figure 1 shows $f_{\rm baryon}(<r_{\rm e})$ versus the total stellar mass in TNG50, color-coded by $r_{\rm e}$: there is a strong dependence of $f_{\rm baryon}$ on stellar mass. Galaxies of low stellar mass ($M_{*}\lesssim 10^{8.5}\,\rm M_{\odot}$) have baryon fractions in the range $f_{\rm baryon}(<r_{\rm e})\sim 0.1-0.3$ and thus are dark matter-dominated. Above $M_{*}\sim 10^{8.5}\,\rm M_{\odot}$, $f_{\rm baryon}(<r_{\rm e})$ increases rapidly in TNG50, and the most massive systems ($M_{*}\sim 10^{10}\,\rm M_{\odot}$) are baryon-dominated. At fixed stellar mass, the scatter in $f_{\rm baryon}(<r_{\rm e})$ among individual galaxies is modest ($\sigma(f_{\rm baryon}|M_{*}=10^{8}\,\rm M_{\odot})\approx 0.05$), and covariant with $r_{\rm e}$, so that larger effective radii imply lower values of $f_{\rm baryon}$.
@@
These simulation results are in remarkably good agreement with the measurements from de Graaff et al. (2024), inferred from JWST observations of low-mass galaxies ($M_{*}\sim 10^{7-9}\,\rm M_{\odot}$) at $z\sim 6-7$. The observational estimates of the baryonic mass were obtained from spectral energy distribution (SED) modeling to low-resolution spectroscopy, while the total mass estimates were inferred dynamically from measurements of ionized gas kinematics. We emphasize that these observed estimates differ conceptually in important ways from those for the simulated baryon fractions. For example, the observed effective radius represents the projected half-light radius of the emission line instead of a half-mass radius, although we find that these are $\sim 1\,\rm kpc$ in both simulated and observed galaxies. The stellar masses were inferred from SED modeling and the cold-gas mass is empirically inferred from the star formation rate (SFR); both observational estimates carry significant systematic uncertainties that are not reflected in the error bars shown.
For a complete discussion of the systematic uncertainties in the different mass estimates we refer the reader to de Graaff et al. 2024, but we provide a brief summary here. The stellar mass in particular can be underestimated by as much as a factor 10 in extreme scenarios where only the integrated light of the galaxy is considered in the SED fitting and a luminous young stellar population outshines the underlying population of older stars in the galaxy (e.g. Giménez-Arteaga et al., 2023). However, the spatially-resolved analysis of one of the galaxies shown in Figure 1, with an SED that is representative of the full $z\sim 6$ sample, yields an integrated stellar mass that is consistent with the mass inferred from SED fitting to the integrated light (Baker et al., 2023). The inferred gas masses are also uncertain, as these were obtained from the conversion of a star formation surface density under the assumption of a constant star formation efficiency (Kennicutt 1998; for further discussion see also Section 5), whereas we may expect this efficiency to increase toward higher redshift (Tacconi et al., 2020). The star formation efficiency in the low-mass, high-redshift regime is still poorly constrained observationally, but we note that Price et al. (2020) found that for galaxies at cosmic noon the gas masses estimated using the empirical relations by Kennicutt (1998) and Tacconi et al. (2020) differ by only 0.13 dex. Such a systematic difference (i.e a factor 1.3 in the inferred $f_{\rm baryon}$) is still consistent within the scatter seen among the individual galaxies in TNG50.
Finally, we note that observations are subject to projection effects that can lead to an underestimated circular velocity. This may for example explain the data point for which the baryonic mass exceeds the total dynamical mass ($M_{\rm baryon}>M_{\rm dyn}$). In light of these different issues, the broad agreement found between the observations and the TNG50 data shown in Figure 1 is striking.
@@
Figure 1: Baryonic mass fraction within the stellar half-mass radius (TNG50) and the emission line half-light radius (JWST; de Graaff et al. 2024) of galaxies vs. total stellar mass. Throughout, by baryonic fraction we mean the ratio between the mass in gas and stars and the total mass of the galaxy (see Eq. 1). The TNG50 simulation predicts low baryon fractions for $M_{*}<10^{9}\,\rm M_{\odot}$ and a sharp increase in $f_{\rm baryon}$ toward higher masses. The half-mass radii are small ($\sim 1\,\rm kpc$ in physical units) and, at fixed stellar mass, are anti-correlated with the baryon fraction. With the exception of one source, which may have high $f_{\rm baryon}$ due to an underestimated dynamical mass, the baryon fractions measured at $z\sim 6$ with JWST are in good agreement with the simulated galaxies.
@@ 4.2 The central mass component budget:
For comparison, we also extract the mass components of central galaxies in the EAGLE cosmological hydrodynamical simulation (Schaye et al., 2015; Crain et al., 2015), specifically the reference model run for a volume of $100^{3}$ comoving Mpc3 (Ref-L100N1504), using halo and subhalo catalogs that were constructed with the exact same methods as for the TNG simulations [^2] Namely, we use a version of EAGLE that has been rewritten and analyzed exactly as the TNG data (see Nelson et al., 2019b, data release paper). This enables us to perform a uniform analysis on all considered simulations.. The EAGLE simulation has similar baryonic and dark matter particle mass resolution to the TNG100 simulation and assumes a similar cosmology. In the context of this paper, the most important difference between the two simulations is the subgrid galaxy formation model, as the two use different prescriptions for, e.g., star formation, stellar feedback and black hole feedback. It is therefore interesting that – just like the TNG simulations – EAGLE predicts $f_{\rm baryon}(<1\,\rm kpc)$ depends strongly on stellar mass at fixed redshift, while $f_{\rm baryon}(<1\,\rm kpc~{}|~{}M_{*})$ is only weakly dependent on redshift, noting that the apparent decrease from $z=1$ to $z=0$ likely reflects the fact that the mass profiles of $z\sim 0$ EAGLE galaxies are affected by the resolution of the simulation for $r\lesssim 2\,\rm kpc$ (see Schaller et al., 2015; de Graaff et al., 2022).
Turning to the descendants of the selected $z=6$ galaxies in TNG50, we find that for any individual descendant $f_{\rm baryon}(<1\,\rm kpc)$ increases rapidly toward lower redshift: while the central region is still dark matter-dominated at $z=6$ with $f_{\rm baryon}(<1\,\rm kpc)\sim 0.25$, it is highly baryon-dominated at $z\sim 2$ $(f_{\rm baryon}(<1\,\rm kpc)\sim 0.7$). Considering the strong dependence of $f_{\rm baryon}$ on $M_{*}$, this evolution can be understood by the rapid growth of total stellar mass over time found in Figure 2. This $M_{*}$-dependence also explains the redshift-dependent ensemble scatter (i.e. the width of the contours): the sample at $z=6$ was selected in a narrow range in stellar mass, where the scatter in $f_{\rm baryon}$ is small (see also Figure 1). As the stellar mass distribution broadens toward lower redshift (Figure 2), the scatter in $f_{\rm baryon}(<1\,\rm kpc)$ at fixed redshift also increases.
@@
Toward lower redshift the stellar-to-gas mass ratio increases sharply for the $z=6$ descendants, and this central build-up of stellar mass coincides with the rapid growth of the total stellar mass (Figure 2), reminiscent of the compaction phase found in zoom-in simulations of massive galaxies at $z\sim 2-4$ (e.g. Zolotov et al., 2015; Lapiner et al., 2023). At $z\lesssim 2$ we find that $M_{\rm gas}(<1\,\rm kpc)\sim 0$, and the mass ratio therefore plateaus at $M_{*}/M_{\rm gas}(\lesssim 1\,\rm kpc)\sim 100$ due to our imposed gas mass floor. This likely reflects the kinetic feedback mode from active galactic nuclei (AGN) in the IllustrisTNG model (Pillepich et al., 2018b), in which black hole-driven winds remove gas from their surroundings.
@@
Putting the results of Figure 3 together, we can gain insight into the physical processes behind the redshift-independence of $f_{\rm baryon}$ at fixed stellar mass. The strong dependence of $f_{\rm baryon}(<1\,\rm kpc)$ on stellar mass implies there is a $M_{*}$-dependent balance between the cooling of baryons on the one hand, and feedback from stellar evolution and/or AGN on the other hand. Indeed, Nelson et al. (2019a) showed that this feedback in TNG50 depends strongly on stellar mass: at $z>4$, where feedback is driven primarily by stellar winds and supernovae, the mass loading factor is a factor $\approx 30$ higher for low-mass ($\sim 10^{8}\,\rm M_{\odot}$) galaxies than for the highest masses ($\sim 10^{11}\,\rm M_{\odot}$), and explains why at these high redshifts the most massive galaxies are able to rapidly form a stellar mass-dominated center.
Toward lower redshifts the mass loading factor increases, especially for more massive galaxies, and the $M_{*}$-dependence for the feedback is thus reduced. Yet, despite this redshift evolution in the feedback efficiency, which implies that gas can be expelled more easily from the centers of galaxies at all masses at lower redshift, $f_{\rm baryon}(<1\,\rm kpc~{}|~{}M_{*})$ remains constant with redshift. This can be partially explained by the fact that the cooling of baryons is also redshift-dependent. However, perhaps more importantly, the ratio of $M_{\rm gas}/M_{*}$ changes with redshift as well: at $z<3$ a substantial fraction of the baryonic mass is in the form of stars (even at low stellar mass), which are not affected explicitly by feedback. We note that AGN feedback, which becomes prominent at $z\lesssim 2$ and high stellar masses in TNG50, therefore also plays little role in the evolution of $f_{\rm baryon}$, as by this time the stellar mass-dominated center has already been formed.
It is thus likely the interplay between cooling, feedback, and central stellar mass growth that conspires to the redshift-independence of $f_{\rm baryon}(<1\,\rm kpc~{}|~{}M_{*})$ in the TNG simulations. The fact that the EAGLE simulation with a very different subgrid model shows a qualitatively similar trend is intriguing. At $z>5$ there appears to be a slight upturn in $f_{\rm baryon}(<1\,\rm kpc~{}|~{}M_{*})$ in the EAGLE simulation, which may reflect physical differences in the stellar feedback or cooling prescriptions with respect to the TNG model at high redshifts.
@@ 4.3 Comparison with observations at cosmic noon
In Figure 4 we show the dark matter fraction, $f_{\rm DM}(<r_{\rm e})\equiv 1-f_{\rm baryon}(<r_{\rm e})$, as a function of redshift. For the TNG50 $z=6$ descendant sample the median evolution in $f_{\rm DM}$ (solid line) is color-coded by the median stellar mass at the given redshift, and contours indicate the variation within the population. Similar to Figure 3, the ensemble median and scatter in $f_{\rm DM}(r_{\rm e})$ at fixed redshift depend strongly on the stellar mass distribution, which broadens in width and shifts to higher stellar masses toward lower redshift. The evolution in $r_{\rm e}$ also plays an important role, leading to a slight upturn in $f_{\rm DM}(<r_{\rm e})$ at $z<1$, and further increasing the ensemble scatter (see also Figure 1). We also show results at $z\sim 1-3$ from various ground-based studies that obtained dynamical masses from ionized gas kinematics (Wuyts et al., 2016; Genzel et al., 2020; Price et al., 2020; Puglisi et al., 2023). All observed data points are color-coded by the total stellar mass, typically inferred from SED modeling to photometric data.
@@ 5 Discussion & Conclusions
We have used the cosmological simulation TNG50 to extract theoretical expectations for the stellar, gas and dark matter mass components in the central parts of high-redshift galaxies. We have explored these components’ fractional contributions as a function of stellar mass and redshift. In particular, we have shown that TNG50 predicts that the central baryon fraction of galaxies is independent of redshift at fixed stellar mass: the centers of low stellar mass galaxies have always been dark matter-dominated. The central baryon fraction, is however, a strong function of stellar mass at all redshifts, and so greatly increases along the evolutionary paths of individual galaxies. The central stellar-to-gas mass ratio, on the other hand, grows strongly towards lower redshifts, even at a fixed stellar mass.
Turning to high redshifts, we compare these results to recent measurements of the mass budget at $z\sim 6$ from JWST observations of galaxies of low stellar mass ($M_{*}\sim 10^{7-9}\,\rm M_{\odot}$; de Graaff et al., 2024). Overall, these observational results appear in good agreement with TNG50 galaxies at $z=6$: the baryonic mass fractions are approximately equally low at low stellar masses $f_{\rm baryon}(<1\,\rm kpc;M_{*}=10^{8-9}\,\rm M_{\odot})\sim 0.2-0.4$. The relative mass fraction among the baryonic components (stars and gas) also agrees at least qualitatively. We find $M_{*}(<1\,\rm kpc)\ll M_{\rm gas}(<1kpc)$ for both the simulated and observed galaxies at $z\sim 6$ and $M_{*}<10^{9}\,\rm M_{\odot}$. Quantitatively, on the other hand, we find that – at face value – the stellar-to-gas mass ratio is a factor two higher in the TNG50 simulated systems compared to the observations. However, we deem this comparison inconclusive as this mass ratio in TNG depends on the resolution of the simulation (at the $20-40\%$ level across a factor of 16 in mass resolution at $z\sim 5-6$ and $M_{*}\sim 10^{8-9}\,\rm M_{\odot}$).
@@
The comparison in the mass budget between the high-redshift observations and simulations demonstrates the great potential constraining power of such observations. The TNG50 simulation makes clear predictions for the central baryon fraction and its dependence on stellar mass, and (lack thereof) on redshift. This is similarly true for the stellar-to-gas mass ratio, which also appears more sensitive to the details of the simulation (resolution, subgrid model). Therefore, systematically mapping these mass ratios as a function of stellar mass and redshift through observations would provide direct insight into galaxy mass assembly. Current surveys with JWST that simultaneously probe the stellar population, ISM and kinematic properties of galaxies at $z>1$ across a wide range in stellar mass provide an ideal dataset for such a study (Maseda et al., 2024; Eisenstein et al., 2023). Combined with a careful modeling analysis of observational biases by the use of mock observations from simulations, these observations will be able to place strong constraints on theoretical models and constrain the fundamental question: how efficient was star formation in the early universe?